A normal working week consists of 5 days of 8 hours per day. The normal rate of pay is RM2.40 

per hour. Overtime rate is calculated at time and a half.

Average output per operative for a 40-hours week = 120 articles.

In order to increase output and eliminate overtime, the company decided to change to a system of 

payment by result start from August 2021. The information on the new system is as follows:

 Time rate: RM2.40 per hour

 Basic time allowed for 15 articles is 5 hours

 Piece work rate: 20% higher than the current time rate per unit.


Calculate the earnings of Alexis if he produces 160 articles in the first week of August 2021.


1
Expert's answer
2021-10-20T15:55:42-0400

Solution:

Basic hrs allowed per week = 5hrs "\times" 5 = 25hrs

Basic time rate = RM2.40

Basic pay for the first week of August 2021 = 25hrs "\times" 2.40 = RM60


Articles produced under the basic pay = 15 "\times" 5 = 75 articles

Extra articles produced = 160 – 75 = 85 articles


Piece work rate = 2.40 "\times"1.20 = RM2.88

Piece work rate pay = 85 "\times"2.88 = RM244.8

Total weekly pay = 60 + 244.8 = RM304.8


The earnings of Alexis if he produces 160 articles in the first week of August 2021 =  RM304.80